ADNAN MUHAMMED

Grade 12,

lim x → 0 (x cosx - log (1+x) )/x 2 is equal to?

lim x → 0  (x cosx - log (1+x) )/x2 is equal to?

Grade:12th Pass

2 Answers

jitender lakhanpal
12 years ago

Dear aditi

lim x → 0  (x cosx - log (1+x) )/x2

cosx = 1 - x2/2! +x4 / 4! ............

x cosx  = x - x3 /2! + x5/4!..........

log (1+x) = x -x2/2 + x4/2...........

so put the above expansion in the limits we get the limit 1/2

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender

santhanagopalan raghavan
12 years ago

simplest way is to use l-hospitals rule.

d^n(numerator)/d^n(denominator) until 0/0 form is removed.then substitute to obtain the limit.

(cosx - xsinx - 1/(1+x))/2x        Now since 0/0 form persists again differentiate

(-sinx - sinx -xcosx + 1/(1+x^2))/2     put x=0 to obtain the limit as,1/2